OpenStudy (anonymous):

when seven times a certain number is subtracted from 15, result is -4. Find the number.

OpenStudy (anonymous):

make this into an equation, do you know how?

OpenStudy (anonymous):

7 x N -15 = -4

OpenStudy (anonymous):

yes! now solve the equation

OpenStudy (anonymous):

x = 11/7

OpenStudy (anonymous):

but for some reason my book says the answer is 19/7

OpenStudy (anonymous):

oh! well it could be a typo because i got x=11/7 too

OpenStudy (anonymous):

okay nevermind i see what went wrong!

OpenStudy (anonymous):

it says when 7 times a number is subtracted from 15 meaning you're doing 15-7n=-4